# Explainability of Text Processing and Retrieval Methods: A Survey

### v2

@@ 6.6. Local explanations: LIME and SHAP inspired approaches
We now turn to studies that propose methods for constructing local explanations. As discussed in Section 4.2, LIME based local, post-hoc explanations have become popular for classification tasks. Following (Ribeiro et al., 2016), Verma and Ganguly (2019) proposed LIRME, an adaptation of LIME that provides pointwise explanations for document ranking. $\mathcal{M}_{C}(Q,D)$, the score that an IR model $\mathcal{M}_{C}$ computes for a query-document pair $(Q,D)$, is explained by a vector of weights corresponding to different tokens, where the weight for a token indicates its contribution to $\mathcal{M}_{C}(Q,D)$. To estimate this explanation vector, perturbed versions of $D$ were generated by selecting subsets of words from $D$, and the effects of the perturbations on the retrieval score assigned by $\mathcal{M}_{C}$ were noted. Experiments were conducted on the TREC-8 adhoc collection. The authors also proposed two metrics to evaluate the generated explanations: consistency and correctness. Consistency measures how stable the explanation vector is across perturbations, while correctness quantifies whether an explanation vector assigns higher weights to “relevant” terms. Perhaps the biggest flaw of LIRME is the overly simplistic definition of ground truth about which terms are actually relevant: this is determined based on the weights of terms occurring in relevant documents, calculated using a simple statistical model (LM-JM).
@@
Singh and Anand (2019) also incorporated ideas from LIME to create pointwise explanations. Their method, EXS, uses a linear SVM as the simple model that approximates an arbitrary, black-box ranker. As usual, perturbed versions of a document $D$ are constructed by deleting random words from random positions in $D$. The authors mainly focus on three different ways to estimate the relevance of a perturbed instance of $D$:  (a) top-k binary: the perturbed version is regarded as relevant if its score, as computed by the simple model (linear SVM), is more than that of the $k^{th}$ document in the original ranked list, and is regarded as non-relevant otherwise;  (b) rank based: this is similar to the above strategy, except that the relevance score for a relevant document decays with its rank, instead of being 1; and  (c) score based: In this method, relevance scores are real-valued, rather than binary and are determined by the relative difference in score between the perturbed document and the top-ranked document. The form of an explanation is similar to that for LIRME: it consists of a graph displaying which words are strong indicators of relevance for a particular query. EXS was used to explain neural retrieval models such as DRMM (Guo et al., 2016) and DESM (Nalisnick et al., 2016), on traditional news collections used at TREC (AP, LATIMES, Robust04 and Financial Times).
@@
Fernando et al. (2019) explored the use of Shapley values (§ 4.3.2) for model-introspective explanations for IR. They used DeepSHAP, which internally computes the Shapley values of terms present in a document (a high value indicates a bigger contribution to the prediction of relevance). The authors compared various approaches for creating a reference document, in order to determine which reference document provides the best explanations for retrieval models. For example, reference documents were constructed by averaging the embeddings of either all words in the vocabulary or a set of words randomly sampled from the collection, excluding the top 1000 documents retrieved by the query. These explanation strategies were investigated on three NRMs: DRMM, MatchPyramid (Pang et al., 2016b), and PACRR-DRMM (McDonald et al., 2018). LIME based explanations were used as a baseline. Although these models use standard neural network architectures, empirical results showed that no single reference document yields the best explanations across different neural ranking models. Further, the performance of these retrieval models is also sensitive with respect to different reference inputs.
@@ 6.7. Extracting and evaluating alignment rationales as local explanations
As discused in Section 3.4.2, a rationale is a set of tokens in an input $I$ that best explains a model’s decision for $I$. In an IR context, alignment rationales, which establish a correspondence between portions (or spans) of a query and a document, are particularly useful. Consider the hypothetical query Where is SIGIR 2022? (Kim et al., 2022), in response to which a neural model may plausibly retrieve the span “SIGIR 2022 will be held in Madrid”. The underlined parts of the query and the document constitute an alignment rationale (AR), since Madrid, a location, is a response to the Where is portion of the query. [^5] Clearly, for conventional IR models, alignment rationales are limited to term matches between the query and the document. Kim et al. (2022) proposed various perturbation-based metrics to evaluate how faithfully ARs reflect a model’s behaviour. Deletions and substitutions were considered as possible perturbations. Empirical findings with a BERT-based ranker (DAI-MAX (Dai and Callan, 2019)) that was trained on the MSMARCO document ranking task suggest that metrics based on substitutions are more reasonable compared to those based on deletions for measuring faithfulness.
Leonhardt et al. (2023) proposed a rationale-based ‘interpretable’ ranking algorithm. The top 20 sentences were extracted from a document as a rationale; a ‘sparse’ representation consisting only of these sentences was used to rank each document. Two sentence-scoring approaches were explored:  (a) an end-to-end gradient-based optimisation technique that chooses sentences in order to maximise retrieval effectiveness in the ranking phase; and  (b) choosing sentences deterministically, e.g., based on their tf-idf scores with respect to the query. For ranking, a monoBERT-style architecture was used with the query and the document rationales. The system’s retrieval effectiveness on the TREC-DL, Core 17, Clueweb09, and BEIR (Thakur et al., 2021) datasets was at par with state-of-the-art models (BERT-3S (Akkalyoncu Yilmaz et al., 2019), Doc-labeled (Dai and Callan, 2019) (DAI-MAXP), BERT-CLS (Nogueira and Cho, 2019) (monoBERT)). Additionally, the quality of the rationales was evaluated by humans.
@@
Pandian et al. (2024) proposed that the quality of rationales be evaluated using intrinsic and extrinsic measures. Intrinsic evaluation quantifies the ‘consistency’ (correlation) between the ranks of the top-$k$ documents when scores are computed using the full-text vs. only rationales (specifically, the concatenation of the top $m$ rationales, where rationale scores were computed using occlusion). [^6] In occlusion (§4.3), document segments are masked in turn; the resultant changes in the document’s score indicate the importance of the respective segment. On the other hand, extrinsic evaluation quantifies how well the top $m$ rationales overlap with the relevant passages of the documents. Relevant passages were identified by aligning the MSMARCO passage and document datasets via the “QnA” version of MSMARCO. Experiments with these datasets suggest that the retrieval effectiveness of a ranking model (both BM25 and NRMs e.g., ColBERT) does not correlate with the proposed intrinsic or extrinsic effectiveness measures.
6.8. Using query expansion for local, listwise explanations
@@
Query expansion (QE) based approaches (see § 3.4.1) approximate a complex retrieval algorithm with a simpler one, by adding ‘explanation’ terms to the original query. IntentEXS (Singh and Anand, 2020) and Multiplex (Lyu and Anand, 2023) are both examples of such approaches, and broadly follow a common framework. A ranked list is viewed as a set of preference pairs of documents. The goal is to identify a set of explanation terms to add to the original query, such that a simple IR model (e.g., BM25) ranks most preference pairs in the same order as a given NRM. While IntentEXS uses a greedy preference pair coverage algorithm to maximise the above objective, Multiplex uses the modern Geno method (Laue et al., 2019) for optimisation.
@@
IntentEXS used LMJM as the simple ranker to explain three pre-BERT neural models: DRMM (Guo et al., 2016), P-DRMM (McDonald et al., 2018), M-Pyramid (Pang et al., 2016a). The effectiveness of the explanation terms was evaluated using two metrics. Fidelity, or predictive accuracy, was measured using the correlation (Kendall’s $\tau$) between the ranked lists of documents retrieved by LMJM and the NRMs. Next, expansion terms generated using RM3 (Lavrenko and Croft, 2001), EMB (Diaz et al., 2016) or DESM (Nalisnick et al., 2016) were regarded as ground truth, and descriptive accuracy was measured in terms of the overlap between the explanation terms and this ‘ground truth’. Experiments were conducted on Clueweb09b (Callan et al., 2009). When the entire ranked list was considered, Kendall’s $\tau$ was found to be <0.5<0.5 for the three NRMs. When only the top $10$ documents were considered, a $\tau$ greater than $0.5$ $(0.537)$ was observed for only P-DRMM. Given these low $\tau$ values, this method does not seem to adequately explain the workings of NRMs. The authors showed, however, that the framework can be used to identify model and temporal biases. For instance, in response to the query ‘fidel castro and electoral college 2008 results’, DRMM considers documents related to events in 2004 or later. Similarly, for the query ‘how to find mean’, DRMM correctly captures the statistical intent behind the query, whereas DESM fails to do so. Multiplex, on the other hand, attempted to explain three NRMs (DRMM, monoBERT and DPR (Karpukhin et al., 2020)) using a combination of three simple rankers: i) term matching (BM25), ii) position aware, and iii) semantic similarity. A fidelity measure (the extent to which preference pairs are in agreement with the original ranking) was the evaluation metric. Not all pairs were considered, however; evaluation was limited to a random sample of preference pairs for which the rank difference was large. Within this restricted setup, Multiplex achieved a high fidelity score on Clueweb09b and TREC-DL 2019.
@@
LIRME (Verma and Ganguly, 2019) Adaptations of LIME TREC-8 github.com/gdebasis/irexplain
EXS (Singh and Anand, 2019) TREC678, Robust github.com/GarfieldLyu/EXS
Fernando et al. (2019) SHAP-based Robust No URL provided
Kim et al. (2022) Evaluating rationale faithfulness MSMARCO docs; TREC-DL 2019 queries github.com/youngwoo-umass/alignment_rationale
Leonhardt et al. (2023) Ranking based on rationales TREC-DL 2019, Clueweb09b, Core 17, BEIR benchmark No URL provided
Pandian et al. (2024) Extracting + evaluating rationales MSMARCO (pass. & docs); TREC-DL 2020 queries github.com/saranpandian/XAIR-evaluation-metric
Singh and Anand (2020) Explanation terms / expanded queries Clueweb09b github.com/GarfieldLyu/EXS
Multiplex (Lyu and Anand, 2023) Clueweb09b, MSMARCO passage and TREC-DL 2019 queries github.com/GarfieldLyu/RankingExplanation
Yu et al. (2022) Diversity-aware listwise explanation Wiki, MIMICS github.com/PxYu/LiEGe-SIGIR2022
BFS (Llordes et al., 2023) Listwise explanation MSMARCO passage and TREC-DL 2019 queries github.com/micllordes/eliBM25
@@
Yu et al. (2022) also consider the problem of providing a listwise explanation of documents. Strictly speaking, however, their objective is not to explain why a given list of documents was retrieved, or ranked in the given order; it would be more appropriate to term their “explanations” as annotations. These annotations are intended to make it easier for a user to consume the information contained in the list of documents. Yu et al. provide two types of annotations. Given a (relatively broad) query involving multiple aspects, the first type of annotation “explains” which aspects of the query are covered by each document in the ranked list. The second type of annotation highlights the novel aspects covered by a particular document, i.e., the aspects that are not already covered by documents ranked above it. The annotations were created by modifying the encoder and decoder components of a Transformer network. Specifically, inter-document and intra-document attention layers were added.
@@
Experiments were conducted on a newly constructed Wikipedia corpus, and the MIMICS (Zamani et al., 2020) dataset. The title of a Wikipedia article was considered as a query. Each section of the article was treated as a unit of retrieval (i.e., as a single document). The heading of any section was treated as the ground-truth / gold-standard explanation that specifies which aspects of the query are covered by the corresponding document. However, as Wikipedia articles are topically organised, most sections cover only a single aspect of a query. Therefore, documents covering multiple, overlapping query aspects were artificially created. Given a Wikipedia article, three sections were randomly sampled from it (without replacement) at a time. Two documents were created from these three sections such that the two documents had one section in common. Thus, from an article containing $n$ sections, $\lfloor n/3\rfloor\times 2$ documents were created. The second dataset (MIMICS) is a search clarification dataset with the search queries sampled from Bing query logs. For each query, the clarifying questions are considered as query aspects.
@@
A group of models from search result explanation tasks, TextRank (Mihalcea and Tarau, ), a topic sensitive PageRank variant of TextRank, LIME based explanations for monoBERT, KeyBERT (Grootendorst, 2020), KeyBERT with maximal marginal relevance (Carbonell and Goldstein, 1998), GenEx (Rahimi et al., 2021), HistGen (Zhang et al., 2019), NMIR (Hashemi et al., 2021) and BART (Lewis et al., 2020) were used as baselines. The quality of explanations was evaluated with BLEU and ROUGE scores. Additionally, BERTScore (Zhang et al., 2020) was computed to measure the semantic similarity between the generated explanations and ground truth explanations. A divergence based metric, Div, was also employed. The proposed architecture outperforms the baselines in terms of the above metrics. 55todo: 5cut down kora jay, too much lekha hoyechge

### v3

@@ 6.6. Local explanations: LIME and SHAP inspired approaches
We now turn to studies that propose methods for constructing local explanations. As discussed in Section 4.2.1, LIME based local, post-hoc explanations have become popular for classification tasks. Following (lime), 10.1145/3331184.3331377 proposed LIRME, an adaptation of LIME that provides pointwise explanations for document ranking. $\mathcal{M}_{C}(Q,D)$, the score that an IR model $\mathcal{M}_{C}$ computes for a query-document pair $(Q,D)$, is explained by a vector of weights corresponding to different tokens, where the weight for a token indicates its contribution to $\mathcal{M}_{C}(Q,D)$. To estimate this explanation vector, perturbed versions of $D$ were generated by selecting subsets of words from $D$, and the effects of the perturbations on the retrieval score assigned by $\mathcal{M}_{C}$ were noted. Experiments were conducted on the TREC-8 adhoc collection. The authors also proposed two metrics to evaluate the generated explanations: consistency and correctness. Consistency measures how stable the explanation vector is across perturbations, while correctness quantifies whether an explanation vector assigns higher weights to “relevant” terms. Perhaps the biggest flaw of LIRME is the overly simplistic definition of ground truth about which terms are actually relevant: this is determined based on the weights of terms occurring in relevant documents, calculated using a simple statistical model (LM-JM).
@@
10.1145/3289600.3290620 also incorporated ideas from LIME to create pointwise explanations. Their method, EXS, uses a linear SVM as the simple model that approximates an arbitrary, black-box ranker. As usual, perturbed versions of a document $D$ are constructed by deleting random words from random positions in $D$. The authors mainly focus on three different ways to estimate the relevance of a perturbed instance of $D$:  (a) top-k binary: the perturbed version is regarded as relevant if its score, as computed by the simple model (linear SVM), is more than that of the $k^{th}$ document in the original ranked list, and is regarded as non-relevant otherwise;  (b) rank based: this is similar to the above strategy, except that the relevance score for a relevant document decays with its rank, instead of being 1; and  (c) score based: In this method, relevance scores are real-valued, rather than binary and are determined by the relative difference in score between the perturbed document and the top-ranked document. The form of an explanation is similar to that for LIRME: it consists of a graph displaying which words are strong indicators of relevance for a particular query. EXS was used to explain neural retrieval models such as DRMM (drmm) and DESM (10.1145/2872518.2889361), on traditional news collections used at TREC (AP, LATIMES, Robust04 and Financial Times). Our study (ir-explain-resource) involving LIRME and EXS suggests that the explanations generated by these approaches are often unstable; even minor perturbations in the document content can lead to substantial changes in the resulting explanations.
@@
10.1145/3331184.3331312 explored the use of Shapley values (Section 4.2.2) for model-introspective explanations for IR. They used DeepSHAP, which internally computes the Shapley values of terms present in a document (a high value indicates a bigger contribution to the prediction of relevance). The authors compared various approaches for creating a reference document, in order to determine which reference document provides the best explanations for retrieval models. For example, reference documents were constructed by averaging the embeddings of either all words in the vocabulary or a set of words randomly sampled from the collection, excluding the top 1000 documents retrieved by the query. These explanation strategies were investigated on three NRMs: DRMM, MatchPyramid (matchpyramid), and PACRR-DRMM (pacrr-drmm). LIME based explanations were used as a baseline. Although these models use standard neural network architectures, empirical results showed that no single reference document yields the best explanations across different neural ranking models. Further, the performance of these retrieval models is also sensitive with respect to different reference inputs.
@@ 6.7. Extracting and evaluating alignment rationales as local explanations
As discused in Section 3.4.2, a rationale is a set of tokens in an input $I$ that best explains a model’s decision for $I$. In an IR context, alignment rationales, which establish a correspondence between portions (or spans) of a query and a document, are particularly useful. Consider the hypothetical query Where is SIGIR 2026? (textual-alignment-sigir22), in response to which a neural model may plausibly retrieve the span “SIGIR 2026 will be held in Melbourne”. The underlined parts of the query and the document constitute an alignment rationale (AR), since Melbourne, a location, is a response to the Where is portion of the query. [^5] Clearly, for conventional IR models, alignment rationales are limited to term matches between the query and the document. textual-alignment-sigir22 proposed various perturbation-based metrics to evaluate how faithfully ARs reflect a model’s behaviour. Deletions and substitutions were considered as possible perturbations. Empirical findings with a BERT-based ranker (DAI-MAX (dai-maxp)) that was trained on the MSMARCO document ranking task suggest that metrics based on substitutions are more reasonable compared to those based on deletions for measuring faithfulness.
extractive_explanation_tois proposed a rationale-based ‘interpretable’ ranking algorithm. The top 20 sentences were extracted from a document as a rationale; a ‘sparse’ representation consisting only of these sentences was used to rank each document. Two sentence-scoring approaches were explored:  (a) an end-to-end gradient-based optimisation technique that chooses sentences in order to maximise retrieval effectiveness in the ranking phase; and  (b) choosing sentences deterministically, e.g., based on their tf-idf scores with respect to the query. For ranking, a monoBERT-style architecture was used with the query and the document rationales. The system’s retrieval effectiveness on the TREC-DL, Core 17, Clueweb09, and BEIR (thakur2021beir) datasets was at par with state-of-the-art models (BERT-3S (bert-3s), Doc-labeled (dai-maxp) (DAI-MAXP), BERT-CLS (DBLP:journals/corr/abs-1901-04085) (monoBERT)). Additionally, the quality of the rationales was evaluated by humans. jacovi-goldberg-2021-aligning criticised this framework in the context of NLP tasks, arguing that the selected rationales may be uninformative from a user’s perspective. We observe a similar phenomenon in retrieval tasks, where the algorithm tends to highlight sentences that humans do not consider important.
@@
dg-ecir-24 proposed that the quality of rationales be evaluated using intrinsic and extrinsic measures. Intrinsic evaluation quantifies the ‘consistency’ (correlation) between the ranks of the top-$k$ documents when scores are computed using the full-text vs. only rationales (specifically, the concatenation of the top $m$ rationales, where rationale scores were computed using occlusion). [^6] In occlusion (Section 4.2.2), document segments are masked in turn; the resultant changes in the document’s score indicate the importance of the respective segment. On the other hand, extrinsic evaluation quantifies how well the top $m$ rationales overlap with the relevant passages of the documents. Relevant passages were identified by aligning the MSMARCO passage and document datasets via the “QnA” version of MSMARCO. Experiments with these datasets suggest that the retrieval effectiveness of a ranking model (both BM25 and NRMs e.g., ColBERT) does not correlate with the proposed intrinsic or extrinsic effectiveness measures.
6.8. Local, listwise explanations: using query expansion and query aspects
@@
Query expansion (QE) based approaches (see Section  3.4.1) approximate a complex retrieval algorithm with a simpler one, by adding ‘explanation’ terms to the original query. IntentEXS (singh-fat-20) and Multiplex (lijun-ecir-23) are both examples of such approaches, and broadly follow a common framework. A ranked list is viewed as a set of preference pairs of documents. The goal is to identify a set of explanation terms to add to the original query, such that a simple IR model (e.g., BM25) ranks most preference pairs in the same order as a given NRM. While IntentEXS uses a greedy preference pair coverage algorithm to maximise the above objective, Multiplex uses the modern Geno method (geno) for optimisation.
@@
IntentEXS used LMJM as the simple ranker to explain three pre-BERT neural models: DRMM (drmm), P-DRMM (pacrr-drmm), M-Pyramid (pang2016study). The effectiveness of the explanation terms was evaluated using two metrics. Fidelity, or predictive accuracy, was measured using the correlation (Kendall’s $\tau$) between the ranked lists of documents retrieved by LMJM and the NRMs. Next, expansion terms generated using RM3 (10.1145/383952.383972), EMB (diaz-etal-2016-query) or DESM (10.1145/2872518.2889361) were regarded as ground truth, and descriptive accuracy was measured in terms of the overlap between the explanation terms and this ‘ground truth’. Experiments were conducted on Clueweb09b (callan2009clueweb09). When the entire ranked list was considered, Kendall’s $\tau$ was found to be <0.5<0.5 for the three NRMs. When only the top $10$ documents were considered, a $\tau$ greater than $0.5$ $(0.537)$ was observed for only P-DRMM. Given these low $\tau$ values, this method does not seem to adequately explain the workings of NRMs. The authors showed, however, that the framework can be used to identify model and temporal biases. For instance, in response to the query ‘fidel castro and electoral college 2008 results’, DRMM considers documents related to events in 2004 or later. Similarly, for the query ‘how to find mean’, DRMM correctly captures the statistical intent behind the query, whereas DESM fails to do so. Multiplex, on the other hand, attempted to explain three NRMs (DRMM, monoBERT and DPR (dpr)) using a combination of three simple rankers: i) term matching (BM25), ii) position aware, and iii) semantic similarity. A fidelity measure (the extent to which preference pairs are in agreement with the original ranking) was the evaluation metric. Not all pairs were considered, however; evaluation was limited to a random sample of preference pairs for which the rank difference was large. Within this restricted setup, Multiplex achieved a high fidelity score on Clueweb09b and TREC-DL 2019.
@@
LIRME (10.1145/3331184.3331377) Adaptations of LIME TREC-8 github.com/gdebasis/irexplain
EXS (10.1145/3289600.3290620) TREC678, Robust github.com/GarfieldLyu/EXS
10.1145/3331184.3331312 SHAP-based Robust No URL provided
textual-alignment-sigir22 Evaluating rationale faithfulness MSMARCO docs; TREC-DL 2019 queries github.com/youngwoo-umass/alignment_rationale
extractive_explanation_tois Ranking based on rationales TREC-DL 2019, Clueweb09b, Core 17, BEIR benchmark No URL provided
dg-ecir-24 Extracting + evaluating rationales MSMARCO (pass. & docs); TREC-DL 2020 queries github.com/saranpandian/XAIR-evaluation-metric
singh-fat-20 Explanation terms / expanded queries Clueweb09b github.com/GarfieldLyu/EXS
Multiplex (lijun-ecir-23) Clueweb09b, MSMARCO passage and TREC-DL 2019 queries github.com/GarfieldLyu/RankingExplanation
listwise-expl-sigir-22 Diversity-aware listwise explanation Wiki, MIMICS github.com/PxYu/LiEGe-SIGIR2022
BFS (llordes-sigir-23) Listwise explanation MSMARCO passage and TREC-DL 2019 queries github.com/micllordes/eliBM25
RankSHAP (rankshap-DBLP:conf/iclr/ChowdhuryZA25) Adaptations of SHAP for listwise ranking MSMARCO (randomly selected 250 Qs), Robust04 No URL provided
RankingSHAP (rankshap-sigir-25) Synthetic, LETOR 4.0 https://github.com/MariaHeuss/RankingShap
@@
listwise-expl-sigir-22 also consider the problem of providing a listwise explanation of documents. Strictly speaking, however, their objective is not to explain why a given list of documents was retrieved, or ranked in the given order; it would be more appropriate to term their “explanations” as annotations. These annotations are intended to make it easier for a user to consume the information contained in the list of documents. listwise-expl-sigir-22 provide two types of annotations. Given a (relatively broad) query involving multiple aspects, the first type of annotation “explains” which aspects of the query are covered by each document in the ranked list. The second type of annotation highlights the novel aspects covered by a particular document, i.e., the aspects that are not already covered by documents ranked above it. The annotations were created by modifying the encoder and decoder components of a Transformer network. Specifically, inter-document and intra-document attention layers were added.
@@
Experiments were conducted on a newly constructed Wikipedia corpus, and the MIMICS (mimics) dataset. The title of a Wikipedia article was considered as a query. Each section of the article was treated as a unit of retrieval (i.e., as a single document). The heading of any section was treated as the ground-truth / gold-standard explanation that specifies which aspects of the query are covered by the corresponding document. However, as Wikipedia articles are topically organised, most sections cover only a single aspect of a query. Therefore, documents covering multiple, overlapping query aspects were artificially created. Given a Wikipedia article, three sections were randomly sampled from it (without replacement) at a time. Two documents were created from these three sections such that the two documents had one section in common. Thus, from an article containing $n$ sections, $\lfloor n/3\rfloor\times 2$ documents were created. The second dataset (MIMICS) is a search clarification dataset with the search queries sampled from Bing query logs. For each query, the clarifying questions are considered as query aspects.
@@
A group of models from search result explanation tasks, TextRank (mihalcea-tarau-2004-textrank), a topic sensitive PageRank variant of TextRank, LIME based explanations for monoBERT, KeyBERT (grootendorst2020keybert), KeyBERT with maximal marginal relevance (mmr-sigir), GenEx (DBLP:journals/corr/abs-2111-01314), HistGen (outline-generation), NMIR (nmir) and BART (lewis-etal-2020-bart) were used as baselines. The quality of explanations was evaluated with BLEU and ROUGE scores. Additionally, BERTScore (Zhang*2020BERTScore:) was computed to measure the semantic similarity between the generated explanations and ground truth explanations. A divergence based metric, Div, was also employed. The proposed architecture outperforms the baselines in terms of the above metrics.